2. One cubic centimeter of a typical cumulus cloud contains 300 spherical liquid water drops, each with a typical radius of 10 um. For a certain spherical cumulus cloud with a radius of 2.0 km, what is the volume of liquid water that it contains? (Hint: first find the ratio of the volume of liquid water to the volume of cloud.)
Answer #1

Volume of spherical liquid drop
3
where R is radius of spherical liquid drop = 10 *10-6 m
On putting the value , we get
u = (3) π(10-5)
v = 4.19*10-15 m3
Now ,
Volume of 300 spherical liquid drop
V = 300v = 300*4.19*10-15 = 1.257*10-12 m3
Now it is given that the
1 Cm3 of cloud contain 300 Spherical liquid drop , therefore
Volume of liquid in 1 cm3 cloud = 1.257*10-12 m3
10-6 m3 of cloud = 1.257*10-12 m3 of liquid
1 m3 of cloud = 1.257*10-6 m3 of liquid
Now the volume of cloud of radius (r) = 2 km = 2000 m
Volume of cloud
V_{c} = left ( rac{4}{3} ight )pi (r)^{3}
= (5) π(2000)
VC = 33.51*109 m3
Now volume of liquid in cloud will be
Vi 33.51 * 10° * 1.257 * 10-0
VL = 42.12*103 m3
hence the volume of liquid in spherical cumulus cloud would be 42.12*103 m3